OpenStudy (zeesbrat3):

A particle moves on a line away from its initial position so that after t hours it is s = 2t2 +3t miles from its initial position. Find the average velocity of the particle over the interval [1, 4].

OpenStudy (zeesbrat3):

@jim_thompson5910

OpenStudy (zeesbrat3):

@ganeshie8

jimthompson5910 (jim_thompson5910):

Position function \[\Large f(t) = 2t^2+3t\] Velocity function (apply derivative) \[\Large v(t) = f \ ' (t) = 4t+3\] A = average velocity \[\Large A = \frac{1}{b-a}\int_{a}^{b}(v(t)dt)\] \[\Large A = \frac{1}{4-1}\int_{1}^{4}(4t+3)dt\] I'll let you finish.